OpenStudy (anonymous):

If you double the velocity of an object, you increase its kinetic energy by a factor of ______.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

kinetic energy=\[\frac{ 1 }{ 2 }*M*V^2\]

OpenStudy (gorv):

if u double the V thn what will happen

OpenStudy (gorv):

\[K.E=\frac{ 1 }{ 2 }*M*V^2\]

OpenStudy (gorv):

\[K.E2=\frac{ 1 }{ 2 }*M*(2V)^2=\frac{ 1 }{ 2 }*M*V^2*2\]

OpenStudy (gorv):

what u think ???

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

I did not get you, should that be \(4\) ??

OpenStudy (gorv):

yeah typin mistake

OpenStudy (gorv):

that will be 4